A website for bed and breakfast inns gets approximately seven visitors per minute.

Suppose the number of website visitors per minute.

a. Compute the probability of no website visitors in a one-minute period. b. Compute the probability of two or more website visitors in a one-minute period. c. Compute the probability of one or more website visitors in a 30-second period. d. Compute the probability of five or more website visitors in a one-minute period.


1
Expert's answer
2021-05-24T01:02:13-0400

Solution:

"\lambda=7" per min"=7\/ 60" per sec

(a) "P(X=0)=e^{-7} \frac{7^0}{0!}=e^{-7}=0.00091"

(b) "P(X\ge2)=1-P(X=0)-P(X=1)-P(X=2)"

"=1-e^{-7} \frac{7^0}{0!}-e^{-7} \frac{7^1}{1!}-e^{-7} \frac{7^2}{2!}
\\=0.97036"

(c) "\lambda" for 30 sec period"=7(30\/60)=3.5"

"P(X\ge1)=1-P(X=0)-P(X=1)
\\=1-e^{-3.5} \frac{3.5^0}{0!}-e^{-3.5} \frac{3.5^1}{1!}
\\=0.8641"

(d) "P(X\ge5)=1-P(X=0)-P(X=1)-P(X=2)-P(X=3)-P(X=4)-P(X=5)"

"=1-e^{-7} \frac{7^0}{0!}-e^{-7} \frac{7^1}{1!}-e^{-7} \frac{7^2}{2!}-e^{-7} \frac{7^3}{3!}-e^{-7} \frac{7^4}{4!}-e^{-7} \frac{7^5}{5!}"

"=0.69929"